# VPGS-SLAM: Voxel-based Progressive 3D Gaussian SLAM in Large-Scale Scenes

### v1

@@ 3.1 Voxel-based Progressive Scene Representation
Table 1: Camera Tracking Performance on Replica dataset replica . We use ATE RMSE (cm) as the metric. Best results are highlighted as first, second, and third.
Methods $0$ 0059 0106 0169 0181 0207 0054 0233 Avg.
@@
NICE-SLAMniceslam 12.3 14.2 7.9 10.9 13.6 6.8 20.9 9.4 13.2
ESLAMeslam 7.7 8.6 7.8 6.5 9.4 6.2 36.5 4.6 10.7
Co-SLAMcoslam 7.2 11.4 9.5 5.9 11.9 7.3 - - -
Point-SLAMpointslam 10.4 7.9 8.8 22.2 14.9 9.7 28.2 6.3 14.4
GO-SLAMgoslam 5.7 7.5 7.0 7.8 6.8 6.7 8.8 4.9 6.8
PLGSLAMplgslam 7.3 8.2 7.4 6.3 9.2 5.8 30.8 4.2 9.9
Loopy-SLAMloopyslam 4.9 7.7 8.5 7.7 10.6 7.9 14.5 5.2 7.7
@@
SplaTAMsplatam 12.8 10.1 17.7 12.1 11.1 7.5 56.8 4.8 16.6
MonoGSmonogs 9.8 32.1 8.9 10.7 21.8 7.9 17.5 12.4 15.2
Gaussian-SLAMgaussianslam 21.2 12.8 13.5 16.3 21.0 14.3 37.1 11.1 18.4
Loop-Splatloopsplat 6.2 7.2 7.4 10.8 8.5 6.7 16.3 4.8 8.4
@@
Table 2: Camera Tracking Performance on ScanNet dataset scannet . We use ATE RMSE (cm) as the metric. We evaluate on eight sequences following the experimental settings of previous methods.
Figure 3: Qualitative comparison between our proposed method and existing SOTA methods: SplaTAM splatam and Loop-Splat loopsplat . We demonstrate RGB image rendering results on the KITTI odometry dataset kitti . Our method shows improved rendering quality compared to these existing methods.
3.3 2D-3D Gaussian Loop Closure
In large-scale environments, pose drift accumulates significantly, posing a major challenge for existing methods. Furthermore, the fusion of multiple submaps also requires alignment to achieve global consistency. We incorporate a novel loop closure method into our framework based on 3D Gaussians to identify pose corrections for past submaps and keyframes. Although some existing GS-SLAM methods loopsplat have incorporated loop closure, we improve the loop closure optimization by combining 2D photometric loss with voxel-based ICP, which allows better adaptation to domain differences between indoor and outdoor environments. Furthermore, we design an online distillation methods for submap fusion upon loop closure detection, enabling the integration of information from different submaps to achieve global consistency.
Loop Detection and Correction In order to detect when the robot revisits the same place, we extract the BEV visual descriptor with a lightweight method BEVPlace++ bevplace++ . We then compute the cosine similarities of the descriptor across different submap $\alpha$, $\beta$. Upon detecting a loop, we construct pose optimization graph with the loop closure constraints. We utilize both 2D rendering loss and voxel based ICP as loop closure constraints to achieve better adaptation to domain differences. We obtain a set of pose corrections $\{R_{i}|t_{i}\}_{M}$ from PGO, where $M$ denotes the correction for submap $M$. Then we update both the camera pose and attributes of 3D Gaussians:
@@
SplaTAM splatam 2.74 4.02 84.61
MonoGS monogs 3.16 4.45 81.52
Gaussian-SLAM gaussianslam 2.53 3.77 84.65
Splat-SLAM splatslam 2.49 3.68 84.79
@@
Table 3: Scene reconstruction performance on Replica dataset replica . We use accuracy [cm], completion [cm] and completion ratio (%) as the metrics for mesh evaluation.
Methods VKITTI2 vkitti2 KITTI kitti
@@
GO-SLAM goslam 19.29 0.73 15.71 0.51
SplaTAM splatam 18.29 0.69 14.68 0.48
MonoGS monogs 18.45 0.70 14.73 0.49
Gaussian-SLAM gaussianslam 18.78 0.69 14.78 0.50
Loop-Splat loopsplat 19.47 0.76 16.77 0.71
@@
Table 4: Scene Reconstruction Performance comparison on KITTI kitti and VKITTI 2 vkitti2 datasets. We use PSNR and SSIM as the metrics.
Method Metrics Rm0 Rm1 Rm2 Off0 Off1 Off2 Off3 Off4 Avg.
@@
Point- SLAMpointslam Depth 0.53 0.22 0.46 0.30 0.57 0.49 0.51 0.46 0.44
@@
Loopy- SLAMloopyslam Depth 0.33 0.21 0.43 0.24 0.47 0.62 0.37 0.26 0.37
@@
Loop- Splatloopsplat Depth 0.39 0.23 0.52 0.32 0.51 0.63 1.09 0.40 0.51
@@
Table 5: Scene Reconstruction Performance on Replica dataset replica .
@@
Suma suma Surfel 2.94 13.85 8.43 0.94 0.43 1.26 0.47 0.54 2.87 2.95 1.37 3.61
Litamin2 litamin2 NDT 5.84 15.93 10.74 0.85 0.77 2.46 0.91 0.65 2.57 2.13 1.05 4.39
FLOAM floam Point Cloud 5.03 3.27 8.64 0.74 0.35 3.43 0.53 0.63 2.57 2.14 1.08 2.84
KISS-ICP kissicp 3.72 10.35 7.86 2.13 0.51 1.37 0.65 0.63 3.64 2.35 1.49 3.47
Mesh-Loam meshloam
@@
Puma puma Mesh 6.64 32.61 18.55 2.25 0.94 3.34 2.41 0.93 6.34 3.94 4.43 8.23
SLAMesh slamesh 5.51 10.93 13.29 0.83 0.33 3.74 0.71 0.83 5.13 1.15 1.17 4.36
Mesh-Loam meshloam 5.37 3.25 7.45 0.54 0.37 1.74 0.38 0.45 3.35 1.74 0.95 2.77
NeRF-Loam nerfloam NeRF 8.64 20.58 7.45 1.79 0.80 5.01 2.47 0.79 4.76 3.47 3.02 5.88
PIN-LO pinslam 5.84 4.37 9.35 0.85 0.19 1.87 0.53 0.55 3.27 2.35 0.97 3.01
PIN-SLAM pinslam 1.18 3.46 2.69 0.80 0.17 0.37 0.48 0.30 2.58 1.34 0.96 1.43
@@
Table 6: Tracking performance comparison on KITTI dataset kitti . The evaluation is conducted on the LiDAR dataset with motion compensated point cloud. We use ATE RMSE (m) as our metric. “Avg." denotes the average results of all sequences. Note that ours is the only 3DGS-based method able to run successfully on all sequences, so we present only our method as a representative of GS-based approaches.
@@
NICE-SLAM niceslam 6.98ms 28.88ms 68.54ms 1.23s 0.30 0.06M 48.48MB 12.0GB
Vox-Fusion voxfusion 7.02ms 23.34ms 350.9ms 0.47s 1.31 0.06M 43.48MB 17.6GB
Point-SLAM pointslam 6.85ms 19.98ms 25.72ms 10.47s 1.33 0.127M 55.42MB 8.5GB
ESLAM eslam 6.85ms 19.98ms 54.80ms 0.29s 2.82 0.003M 27.12MB 17.3GB
Co-SLAM coslam 6.38ms 14.25ms 63.93ms 0.15s 3.68 0.013M 24.85MB 16.7GB
Loopy-SLAM loopyslam 55.53ms 10.24ms 1.11s 4.08s 2.12 0.127M 60.98MB 13.3GB
@@
SplaTAM splatam 24.23ms 22.83ms 2.18s 1.37s 175.64 0M 273.09MB 18.5GB
Loop-Splat loopsplat 10.28ms 17.53ms 1.03s 1.05s 315.48 0M 93.87MB 7.4GB
@@
Table 7: Runtime and Memory Usage on Replica Room 0. Per-frame runtime is calculated as the total optimization time divided by the sequence length. “-" is because our method does not use iterative tracking, making it infeasible to calculate per-iteration time. The memory usage represents the total memory of the map representation. “Decoder Param." denotes decoder parameters. Note that implicit field-based methods require additional space for their decoders.
4 Experiments
Datasets. We evaluate VPGS-SLAM on a variety of scenes from different datasets: a) Replica Dataset replica . 8 small room scenes (approximately $6.5m\times 4.2m\times 2.7m$ ). We use this dataset to evaluate the reconstruction and localization accuracy in small-scale environments. b) ScanNet dataset scannet . Real-world scenes with long sequences (more than 5000 images) and large-scale indoor scenarios (approximately $7.5m\times 6.6m\times 3.5m$). We use this dataset for large-scale real-world indoor environments. c) KITTI dataset kitti and VKITTI 2 dataset vkitti2 . Urban scenes dataset with long sequences (approximately $500m\times 400m$). We use these datasets to validate the effectiveness of our method in indoor and real-world outdoor city-scale scenes.
Implementation Details We run our system on a desktop PC with NVIDIA RTX 3090 GPU. We set the voxelized corresponding parameter $k=10$. All the MLPs employed in our approach are 2-layer MLPs with ReLU activation; the dimensions of the hidden units are all 32. We set the submap initialization parameters with a distance threshold $d=0.5m$ in indoor scenes and $d=10m$ for outdoor scenes, and a rotation threshold $\omega=50$ degree. The two loss weight $\lambda_{SSIM}$ and $\lambda_{vol}$ are set to 0.2 and 0.001 in our experiments. We use DepthLab depthlab for dense depth image in outdoor scene. For further details of our implementation, please refer to the supplementary.
Scene Reconstruction We report the scene reconstruction results on Replica replica in Tab. 5. The best results are highlighted as first, second, and third. Our method outperforms all 3DGS-based methods with superior rendering performance, demonstrating its effectiveness in reconstructing small room scenes. We also evaluate our method on the KITTI dataset in Tab. 5 to verify its effectiveness in large outdoor scenes. In Fig. 3, we visualize the rendered results and compare them with SplaTAM and MonoGS, demonstrating our clearly superior results.
[…]
Memory and Time analysis In Tab. 7, we present the runtime, memory usage, and Peak GPU memory of our method and other SOTA methods. In contrast to GS-based methods, our method employs a progressive map representation with multiple submaps, reducing excessive online GPU usage. This approach reduces GPU requirements by a factor of 2-3 compared to SplaTAM splatam , enabling our algorithm to be effectively utilized on edge computing platforms. Our voxelized scene representation method also significantly improves map representation efficiency, reducing the required memory six times compared to SplaTAM splatam .
Method Replica replica KITTI kitti
@@ 4 Experiments
Table 8: Ablation study on the Replica replica and KITTI kitti dataset. “H" denotes hours. “-" denotes fail. The full model demonstrates superior pose estimation accuracy while maintaining faster training/rendering speed and lower memory consumption.
4.1 Ablation Study
@@
In this section, we validate the effectiveness of each module in our algorithm in both indoor and outdoor datasets, shown in Tab. 8. First, we explore the impact of voxelized scene representation and find that this module significantly enhances mapping efficiency in dense scenes. Next, we compare the effectiveness of progressive mapping with multiple submaps, which considerably reduces GPU computation requirements while improving tracking accuracy. This multi-map representation confines pose errors within each submap, effectively mitigating cumulative drift. We then assess the effectiveness of 2D-3D fusion tracking. Our findings indicate that 2D photometric loss performs better in indoor and simulated environments, whereas 3D geometric cues are more beneficial for outdoor scene tracking. Compared with 3D gaussian-based tracking, our tracking method can achieve more accurate matching with 2D-3D information fusion. Finally, we validate the loop closure module’s effectiveness, showing that it significantly reduces cumulative pose errors.
5 Conclusion

### v2

@@ III-A Voxel-based Progressive Scene Representation
TABLE I: Camera Tracking Performance on Replica dataset [45]. We use ATE RMSE (cm) as the metric. Best results are highlighted as first, second, and third.
Methods 0000 0059 0106 0169 0181 0207 0054 0233 Avg.
@@
NICE-SLAM[68] 12.3 14.2 7.9 10.9 13.6 6.8 20.9 9.4 13.2
ESLAM[20] 7.7 8.6 7.8 6.5 9.4 6.2 36.5 4.6 10.7
Co-SLAM[51] 7.2 11.4 9.5 5.9 11.9 7.3 - - -
Point-SLAM[42] 10.4 7.9 8.8 22.2 14.9 9.7 28.2 6.3 14.4
GO-SLAM[63] 5.7 7.5 7.0 7.8 6.8 6.7 8.8 4.9 6.8
PLGSLAM[14] 7.3 8.2 7.4 6.3 9.2 5.8 30.8 4.2 9.9
Loopy-SLAM[26] 4.9 7.7 8.5 7.7 10.6 7.9 14.5 5.2 7.7
@@
SplaTAM[21] 12.8 10.1 17.7 12.1 11.1 7.5 56.8 4.8 16.6
MonoGS[31] 9.8 32.1 8.9 10.7 21.8 7.9 17.5 12.4 15.2
Gaussian-SLAM[60] 21.2 12.8 13.5 16.3 21.0 14.3 37.1 11.1 18.4
Loop-Splat[65] 6.2 7.2 7.4 10.8 8.5 6.7 16.3 4.8 8.4
@@
TABLE II: Camera Tracking Performance on ScanNet dataset [8]. We use ATE RMSE (cm) as the metric. We evaluate on eight sequences following the experimental settings of previous methods.
Figure 4: Qualitative comparison between our proposed method and existing SOTA methods: SplaTAM [21] and Loop-Splat [65]. We demonstrate RGB image rendering results on the KITTI odometry dataset [17]. Our method shows improved rendering quality compared to these existing methods.
Figure 5: Visualization of RGB image rendering results on the KITTI odometry dataset sequence 02. Qualitative comparison between our proposed method and existing SOTA methods: SplaTAM [21] and Loop-Splat [65]. We demonstrate RGB image rendering results on the KITTI odometry dataset. Our method shows improved rendering quality compared to these existing methods.
Figure 6: Visualization of RGB image rendering results on the KITTI odometry dataset sequence 03. Qualitative comparison between our proposed method and existing SOTA methods: SplaTAM [21] and Loop-Splat [65]. We demonstrate RGB image rendering results on the KITTI odometry dataset. Our method shows improved rendering quality compared to these existing methods.
Figure 7: Mesh reconstruction results on Newer College dataset [40].
III-C 2D-3D Gaussian Loop Closure
In large-scale environments, pose drift accumulates significantly, posing a major challenge for existing methods. Furthermore, the fusion of multiple submaps also requires alignment to achieve global consistency. We incorporate a novel loop closure method into our framework based on 3D Gaussians to identify pose corrections for past submaps and keyframes. Although some existing GS-SLAM methods [65] have incorporated loop closure, we improve the loop closure optimization by combining 2D photometric loss with voxel-based ICP, which allows better adaptation to domain differences between indoor and outdoor environments. Furthermore, we design an online distillation methods for submap fusion upon loop closure detection, enabling the integration of information from different submaps to achieve global consistency.
Loop Detection and Correction In order to detect when the robot revisits the same place, we extract the visual descriptor with a lightweight method Netvlad [1]. We then compute the cosine similarities of the descriptor across different submap $\alpha$, $\beta$. Upon detecting a loop, we construct pose optimization graph with the loop closure constraints. We utilize both 2D rendering loss and voxel based ICP as loop closure constraints to achieve better adaptation to domain differences. We obtain a set of pose corrections $\{R_{i}|t_{i}\}_{M}$ from PGO, where $M$ denotes the correction for submap $M$. Then we update both the camera pose and attributes of 3D Gaussians:
@@
SplaTAM [21] 2.74 4.02 84.61
MonoGS [31] 3.16 4.45 81.52
Gaussian-SLAM [60] 2.53 3.77 84.65
Splat-SLAM [43] 2.49 3.68 84.79
@@
TABLE III: Scene Reconstruction Performance on Replica dataset [45]. We use accuracy [cm], completion [cm] and completion ratio (%) as the metrics for mesh evaluation.
Figure 8: Extensive Experiments on KITTI datasets [17]. Our system is a large-scale SLAM framework, with voxel-based progressive 3D Gaussian representation, 2D-3D fusion camera tracking, and 3D Gaussian loop closure. Our framework takes color images and 3D point clouds as input. Our method can achieve accurate and efficient scene reconstruction, camera tracking, and global map generation.
Methods VKITTI2 [3] KITTI [17]
@@
GO-SLAM [63] 19.29 0.73 15.71 0.51
SplaTAM [21] 18.29 0.69 14.68 0.48
MonoGS [31] 18.45 0.70 14.73 0.49
Gaussian-SLAM [60] 18.78 0.69 14.78 0.50
Loop-Splat [65] 19.47 0.76 16.77 0.71
@@
TABLE IV: Scene Reconstruction Performance comparison on KITTI [17] and VKITTI 2 [3] datasets. We use PSNR and SSIM as the metrics.
Method Metrics Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4 Avgerage
@@
Point- SLAM[42] Depth 0.53 0.22 0.46 0.30 0.57 0.49 0.51 0.46 0.44
@@
Loopy- SLAM[26] Depth 0.33 0.21 0.43 0.24 0.47 0.62 0.37 0.26 0.37
@@
Loop- Splat[65] Depth 0.39 0.23 0.52 0.32 0.51 0.63 1.09 0.40 0.51
@@
TABLE V: Scene Reconstruction Performance on Replica dataset [45]. We use depth L1, PSNR, SSIM as our metics.
@@
Suma [2] Surfel 2.94 13.85 8.43 0.94 0.43 1.26 0.47 0.54 2.87 2.95 1.37 3.61
Litamin2 [59] NDT 5.84 15.93 10.74 0.85 0.77 2.46 0.91 0.65 2.57 2.13 1.05 4.39
FLOAM [50] Point Cloud 5.03 3.27 8.64 0.74 0.35 3.43 0.53 0.63 2.57 2.14 1.08 2.84
KISS-ICP [49] 3.72 10.35 7.86 2.13 0.51 1.37 0.65 0.63 3.64 2.35 1.49 3.47
Mesh-Loam [67]
@@
Puma [48] Mesh 6.64 32.61 18.55 2.25 0.94 3.34 2.41 0.93 6.34 3.94 4.43 8.23
SLAMesh [41] 5.51 10.93 13.29 0.83 0.33 3.74 0.71 0.83 5.13 1.15 1.17 4.36
Mesh-Loam [67] 5.37 3.25 7.45 0.54 0.37 1.74 0.38 0.45 3.35 1.74 0.95 2.77
NeRF-Loam [10] NeRF 8.64 20.58 7.45 1.79 0.80 5.01 2.47 0.79 4.76 3.47 3.02 5.88
PIN-LO [36] 5.84 4.37 9.35 0.85 0.19 1.87 0.53 0.55 3.27 2.35 0.97 3.01
PIN-SLAM [36] 1.18 3.46 2.69 0.80 0.17 0.37 0.48 0.30 2.58 1.34 0.96 1.43
@@
TABLE VI: Tracking performance comparison on KITTI dataset [17]. The evaluation is conducted on the LiDAR dataset with motion compensated point cloud. We use ATE RMSE (m) as our metric. “Avg." denotes the average results of all sequences. Note that ours is the only 3DGS-based method able to run successfully on all sequences, so we present only our method as a representative of GS-based approaches.
@@
NICE-SLAM [68] 6.98ms 28.88ms 68.54ms 1.23s 0.30 0.06M 48.48MB 12.0GB
Vox-Fusion [58] 7.02ms 23.34ms 350.9ms 0.47s 1.31 0.06M 43.48MB 17.6GB
Point-SLAM [42] 6.85ms 19.98ms 25.72ms 10.47s 1.33 0.127M 55.42MB 8.5GB
ESLAM [20] 6.85ms 19.98ms 54.80ms 0.29s 2.82 0.003M 27.12MB 17.3GB
Co-SLAM [51] 6.38ms 14.25ms 63.93ms 0.15s 3.68 0.013M 24.85MB 16.7GB
Loopy-SLAM [26] 55.53ms 10.24ms 1.11s 4.08s 2.12 0.127M 60.98MB 13.3GB
@@
SplaTAM [21] 24.23ms 22.83ms 2.18s 1.37s 175.64 0M 273.09MB 18.5GB
Loop-Splat [65] 10.28ms 17.53ms 1.03s 1.05s 315.48 0M 93.87MB 7.4GB
@@
TABLE VII: Runtime and Memory Usage on Replica Room 0. Per-frame runtime is calculated as the total optimization time divided by the sequence length. “-" is because our method does not use iterative tracking, making it infeasible to calculate per-iteration time. The memory usage represents the total memory of the map representation. “Decoder Param." denotes decoder parameters. Note that implicit field-based methods require additional space for their decoders.
Method Replica [45] KITTI [17]
@@
TABLE VIII: Ablation study on the Replica [45] and KITTI [17] dataset. “H" denotes hours. “-" denotes fail. The full model demonstrates superior pose estimation accuracy while maintaining faster training/rendering speed and lower memory consumption.
IV Experiments
Datasets. We evaluate VPGS-SLAM on a variety of scenes from different datasets:
  • Replica Dataset [45]. 8 small room scenes (approximately $6.5m\times 4.2m\times 2.7m$ ). We use this dataset to evaluate the reconstruction and localization accuracy in small-scale environments.
  • ScanNet dataset [8]. Real-world scenes with long sequences (more than 5000 images) and large-scale indoor scenarios (approximately $7.5m\times 6.6m\times 3.5m$). We use this dataset for large-scale real-world indoor environments.
  • KITTI dataset [17] and VKITTI 2 dataset [3]. Urban scenes dataset with long sequences (approximately $500m\times 400m$). We use these datasets to validate the effectiveness of our method in virtual and real-world city-scale scenes.
  • Newer College datasets [40]. This dataset provides diverse trajectories and complex scenes, making it well-suited for assessing large-scale and high-fidelity reconstruction performance.
We use these datasets to validate the effectiveness of our method in indoor and real-world outdoor city-scale scenes.
Implementation Details We run our system on a desktop PC with NVIDIA RTX 3090 GPU. We set the voxelized corresponding parameter $k=10$. All the MLPs employed in our approach are 2-layer MLPs with ReLU activation; the dimensions of the hidden units are all 32. We set the submap initialization parameters with a distance threshold $d=0.5m$ in indoor scenes and $d=10m$ for outdoor scenes, and a rotation threshold $\omega=50$ degree. The two loss weight $\lambda_{SSIM}$ and $\lambda_{vol}$ are set to 0.2 and 0.001 in our experiments. We use DepthLab [29] for dense depth image in outdoor scene.
Scene Reconstruction To evaluate the performance of scene reconstruction, we utilized three datasets: Replica for small room scenes, KITTI and vKITTI for large-scale urban scenes, and Newer College for campus environments. We report the scene reconstruction results on Replica [45] in Tab. V. The best results are highlighted as first, second, and third. Our method outperforms all 3DGS-based methods with superior rendering performance, demonstrating its effectiveness in reconstructing small room scenes. We also visualize the rendering performance in the Replica scene in Fig. 3. To evaluate the effectiveness of our method in large-scale urban scene reconstruction, we test it on the KITTI [17] and VKITTI2 [3] dataset, as shown in Tab. IV. In Fig. 4, Fig. 5, and Fig. 6, we visualize the rendered results in different KITTI sequence and compare them with SplaTAM [21] and MonoGS [31], demonstrating our clearly superior performance on scene reconstruction and rendering. We also visualize the mesh reconstruction results on the Newer College dataset in Fig. 1. The left image shows the reconstruction of the entire sequence, and the right image shows a zoomed-in view of a local area. As can be seen, in addition to the rendering performance, our mesh reconstruction also achieves excellent results.
[…]
Memory and Time analysis In Tab. VII, we present the runtime, memory usage, and Peak GPU memory of our method and other SOTA methods. In contrast to GS-based methods, our method employs a progressive map representation with multiple submaps, reducing excessive online GPU usage. This approach reduces GPU requirements by a factor of 2-3 compared to SplaTAM [21], enabling our algorithm to be effectively utilized on edge computing platforms. Our voxelized scene representation method also significantly improves map representation efficiency, reducing the required memory six times compared to SplaTAM [21].
IV-A Ablation Study
@@
In this section, we validate the effectiveness of each module in our algorithm in both indoor and outdoor datasets, shown in Tab. VIII. First, we explore the impact of voxelized scene representation and find that this module significantly enhances mapping efficiency in dense scenes. Next, we compare the effectiveness of progressive mapping with multiple submaps, which considerably reduces GPU computation requirements while improving tracking accuracy. This multi-map representation confines pose errors within each submap, effectively mitigating cumulative drift. We then assess the effectiveness of 2D-3D fusion tracking. Our findings indicate that 2D photometric loss performs better in indoor and simulated environments, whereas 3D geometric cues are more beneficial for outdoor scene tracking. Compared with 3D gaussian-based tracking, our tracking method can achieve more accurate matching with 2D-3D information fusion. Finally, we validate the loop closure module’s effectiveness, showing that it significantly reduces cumulative pose errors.
V Conclusion